INAF queue observing, 2015 Feb 10-11

Observer: F. Cusano, R. Carini
Telescope Operator: S. Allanson
Support Astronomer: Olga Kuhn
Instrument: LUCI/MODS-LBC

Summary

At the sunset the sky is clear.

Data Summary

Project_RUN Objects Filter ExpTime Notes
skyflat LBC r-sloan,i-sloan    
Piconcelli A1 0745+4734 Ks 800 s COMPLETED
telluric HIP35787_Intermediate Ks    
Piconcelli A2 0801+5210 Ks 40 m COMPLETED
telluric HIP41798_Intermediate Ks    
Piconcelli A3 0900+4215 Ks 800 s COMPLETED
Saracco B RDCS-2 red grating 1 h 30 m  
Sommariva NGC 4535 dual grating 45 m COMPLETED
standard feige67 red and dual grating    
Mignoli MOS 22 SDSSJ1148 red grating 2 h variable seeing COMPLETED
Mignoli LBC_MOS22_red SDSSJ1148 filter r-SLOAN 1 h  
Mignoli LBC_MOS22_red SDSSJ1148 filter Y 42 m  
phot standard S104 filters r-sloan i-sloan z-sloan Y-fan  
sky flat LBC z-sloan i-sloan    

Details

12:12 pm we start calibration for MODS
  • Bias mods1b.20150211.0001-0010 , mods1r.20150211.0001-0010
  • slitless Flat red grating mods1r.0011-0015
  • slitless Flat dual grating mods1b.0011-0015, mods1r.0016-0020 stopped because TO opens the dome.

01:11 we start sky flat for LBC red
  • r-sloan filter
  • i-sloan filter, last files with low counts
  • 01:47 we finish the sky flats

01:57 test images with LUCI

02:20 we start the Piconcelli A1 program, we acquire the target 0745+4734
  • seeing variable due to collimation.
  • executeLUCIscript.sh A1_LS_Ks.acq luci.20150211.0003-0006

02:39 executeLUCIscript.sh A1_LS_Ks.spec luci.0007-0010
  • During the acquisition we note a strip of bad columns in the right lower part od the CCD.
    This in not affecting the spectra because the spectral range does not reach the bad columns.
    The night assistent suggests to not do other programs with LUCI, especially faint ones.
  • the seeing is about 0.9

02:58 we slew to telluric star
  • * executeLUCIscript.sh A1_LS_JHeI_tell.acq luci.0011-0014

03:07 executeLUCIscript.sh A1_LS_JHeI_tell.spec luci.0015-0018

03:11 we slew to Piconcelli A2 program
  • executeLUCIscript.sh A2_LS_Ks.acq luci.20150211.0019-0022
  • variable seeing: 0.8-1.3
  • the compass of the finding chart was wrong.

03:27 starting science. executeLUCIscript.sh A2_LS_Ks.spec luci.20150211.0023-0028
  • 03:48 variable seeing with maximum at 2.0

04:03 we take two spectra more dure the bad variable seeing luci.20150211.0029-0030

04:15 acquisition of telluric star for Piconcelli A3 program
  • executeLUCIscript.sh A3_LS_HeIJ_tell.acq luci.20150211.0031-0032 these spectra are not good: instead of sending acquisition we send .spec
  • executeLUCIscript.sh A3_LS_HeIJ_tell.acq luci.20150211.0033-0035
  • 4:20 seeing 1.0

04:23 executeLUCIscript.sh A3_LS_HeIJ_tell.spec luci.20150211.0036-0037

04:26 acquisition Piconcelli A3
  • executeLUCIscript.sh A3_LS_Ks.acq luci.20150211.0038-0041

04:38 executeLUCIscript.sh A3_LS_Ks.spec luci.20150211.0042-0045
  • seeing about 1.0

04:54 switch to MODS

05:12 slew to Saracco B
  • acqMods mods.7.RDCSm1.acq mods1r.0021-0023
  • seeing 1.2

05:28 execMODS mods.7.RDCSm1_6.obs
  • mods1r.0024-0027
  • seeing 1
  • 5:57 seeing 1.2
  • 6:25 seeing 1.2

06:40 unwrap the telescope, we send the acq again
  • acqMods mods.7.rec.acq mods1r.0028

06:49 execMODS mods.7.RDCSm1_7.obs
  • seeing 0.9
  • 07:20 seeing 2.0
  • seeing is too bad , we stop mods1r.0029-0030

07:32 seeing is decreasing to 1.3, start again the science
  • execMODS mods.7.RDCSm1_7.obs
  • 07:36 seeing it is too bad to go on with Saracco program, we move to Sommariva program

07:45 seeing 3
  • acqMODS ngc4579_mods.acq mods1r.0031-0033
  • execMODS ngc4579_mods.obs
  • There are problems with the collimation, the guide star is too near to the limit of the patrol field, we have to stop. We change the star guide.

08:04 acqMODS ngc4579_mods.acq
  • tecnical problem in the preset
  • we try again the acquisition, but the problems persist, we change target.

08.19 acqMODS ngc4535_mods.acq mods1r.0034-0036

08:25 expecMODS ngc4535_mods.obs mods1r.0037, mods1b.0016
  • seeing is about 1.7-2.0
  • Marco Stangalini is running some test on LBC-r (remote from Rome)

09:18 slew to standard star of Sommariva, acqMods feige67.acq mods1r.0038-39

09:25 execMODS feige67.obs
  • dual grating mods1r.0040-42, mods1b.0017-0019
  • red grating mods1r.0043-45

09:52 seeing is improving, 1.2 we slew to Mignoli MOS22
  • acqMODS mods.22.acq mods1r.004648-==
  • seeing increases again 1.8, it is very variable

10:00 dofpia, /X2, /first, /redonly

10:09 expecMODS mods.22.5.obs mods1r.0049-52
  • run 7_MOS22_r.ob
  • the seeing is about 1.6
  • lbcr.20150211.100927.fits is not good, has been done during the offset of MODS.

10:23 seeing is 1.2
  • 10:36 seeing 1.1
  • 10:45 seeing 1.3

10:47 dofpia, /X2, /redonly

10:52 run 8_MOS22_Y.ob
  • seeing 1.45
  • 11:15 seeing 1.7
  • 11:20 seeing 2.0

11:24 seeing 1.0. execMODS mods.22.6.obs mods1r.0053-556

11:35 dofpia, /X2, /redonly

11:39 run 9_MOS22_r.ob
  • seeing 1.15

12.21 dofpia, /X2, /redonly

12.27 run 10_MOS22_Y.ob.
  • seeing 1.4 * taken only two images

12:38 slew to photometric standard S104 * filter i-sloan, r-sloan,z-sloan, Y-fan

12:58 slew to Mignoli photometry in Y filter
  • run 10_MOS22_Y.ob, other two images

1.15 sky flat in z-sloan and i-sloan (only one image is good)

Comments

 
Topic revision: r12 - 19 May 2015, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback